### v1

# MOGO: Residual Quantized Hierarchical Causal Transformer for High-Quality and Real-Time 3D Human Motion Generation

Abstract
Recent advances in transformer-based text-to-motion generation have led to impressive progress in synthesizing high-quality human motion. Nevertheless, jointly achieving high fidelity, streaming capability, real-time responsiveness, and scalability remains a fundamental challenge. In this paper, we propose MOGO (Motion Generation with One-pass), a novel autoregressive framework tailored for efficient and real-time 3D motion generation. MOGO comprises two key components: (1) MoSA-VQ, a motion scale-adaptive residual vector quantization module that hierarchically discretizes motion sequences with learnable scaling to produce compact yet expressive representations; and (2) RQHC-Transformer, a residual quantized hierarchical causal transformer that generates multi-layer motion tokens in a single forward pass, significantly reducing inference latency. To enhance semantic fidelity, we further introduce a text condition alignment mechanism that improves motion decoding under textual control. Extensive experiments on benchmark datasets including HumanML3D, KIT-ML, and CMP demonstrate that MOGO achieves competitive or superior generation quality compared to state-of-the-art transformer-based methods, while offering substantial improvements in real-time performance, streaming generation, and generalization under zero-shot settings.
1 Introduction
Text-to-motion generation is a rapidly evolving research area with growing importance in virtual environments, such as gaming, AR/VR, and humanoid robotics [1]. Recent advances have leveraged large language model (LLM)-based techniques to generate high-quality 3D human motion from text descriptions, typically using vector-quantized variational autoencoders (VQ-VAE) and autoregressive decoding strategies [2, 3, 4, 5]. Meanwhile, diffusion-based approaches have also demonstrated strong generation performance, particularly in terms of motion fidelity and diversity [6].
Despite these achievements, both diffusion- and LLM-based frameworks face practical limitations [7, 8, 9, 10]. Diffusion models often rely on iterative refinement processes, which introduce significant inference latency and hinder their suitability for real-time or interactive applications [7]. On the other hand, LLM-based models, although autoregressive, typically involve large parameter sizes and long context dependencies, leading to high memory and computation costs that challenge deployment on lightweight or streaming scenarios [9]. A broader discussion of related research trends is provided in Section 4.
To address these issues, we propose MOGO (Motion Generation with One-pass), a transformer-based autoregressive framework that directly predicts motion tokens in a single forward pass. Compared to diffusion-based models, MOGO enables streamable motion decoding without iterative sampling, significantly reducing inference time. Furthermore, unlike LLM-based motion generators, MOGO adopts a lightweight yet expressive transformer architecture with considerably fewer parameters—allowing for real-time token generation while maintaining high motion fidelity.
MOGO is composed of two key components: (1) MoSA-VQ (Motion Scale-Adaptive Residual VQ-VAE), which discretizes motion sequences via hierarchical quantization and learnable scaling, producing compact and adaptive representations [11]; and (2) RQHC-Transformer (Residual Quantized Hierarchical Causal Transformer), which autoregressively generates multi-layer motion tokens in a single pass [12]. As shown in Figure 1, this design allows for the synthesis of coherent and extended motion sequences while preserving streaming capability and generalization across diverse prompts and datasets.
Extensive experiments demonstrate that MOGO achieves superior results on both in-distribution and out-of-distribution scenarios, outperforming state-of-the-art LLM-based motion generators in quality, inference speed, and scalability—making it a robust and real-time-capable solution for practical text-to-motion generation.
Our main contributions:
MOGO: A unified transformer-based framework for one-pass high-fidelity motion generation. It achieves generation quality comparable to state-of-the-art LLM- and diffusion-based methods, while significantly improving real-time efficiency and enabling streamable motion synthesis.
MoSA-VQ: A motion discretization module that employs hierarchical quantization and learnable feature scaling for compact and expressive representations.
RQHC-Transformer: A lightweight hierarchical causal transformer that autoregressively generates multi-layer motion tokens in one forward pass.
All code and demonstration pages will be made publicly available on our GitHub repository upon acceptance.
2 Methods
We introduce MOGO, a unified framework designed for efficient and expressive text-to-motion generation. Mogo could generate a high-quality motion sequence from a textual description, where each frame represents a -dimensional human pose. As shown in Figure 2, our MOGO framework consists of two core components: (1) MoSA-VQ (left) that discretizes motion into multi-level tokens with high precision, and (2) RQHC-Transformer (right) leveraging hierarchical causal forward computation to generate motion tokens collectively in a streamlined manner (algorithms are shown in Appendix C). To enhance zero/few-shot performance, we incorporate text condition alignment (TCA) for improved motion decoding.
2.1 MoSA-VQ: Motion Scale-Adaptive Residual VQ-VAE
To effectively transform continuous human motion sequences into discrete representations suitable for autoregressive modeling, we adopt a residual vector quantized variational autoencoder (RVQ-VAE) architecture. This module, termed MoSA-VQ, serves as the foundation for learning compact yet expressive motion codes. Residual quantization enables progressive refinement of motion features across multiple levels, capturing both coarse and fine-grained temporal dynamics essential for high-fidelity generation.
However, previous RVQ-based approaches [2, 5] often rely on manually calibrated feature ranges, which limit their adaptability and quantization efficiency in complex motion settings. To address this, we introduce a novel learnable feature scaling mechanism, allowing the quantization space to dynamically adapt at each hierarchy level. This design not only eliminates the need for manual tuning but also improves reconstruction accuracy and codebook utilization.
Hierarchical Quantization with Feature Scaling.
Motion encoder outputs a latent representation , where is the number of motion tokens and is the feature dimension. At each level , we perform feature scaling before residual quantization. Let be the residual feature at level 0. We first compute a scaled version of the residual feature:
where and are learnable scale and bias parameters. The input is the residual feature passed from the previous quantization level and represents the information that has not yet been captured.
We then apply vector quantization to the scaled residual:
where maps each input to the nearest codebook entry. This quantized output exists in the scaled feature space, so we invert the transformation to project it back into the original residual space:
The recovered quantized vector captures the portion of the motion representation explained at level , and we update the residual for the next layer by subtracting this contribution:
This process repeats hierarchically across all levels, resulting in a set of quantized vectors . The final quantized latent representation is reconstructed by summing contributions from all layers:
which is subsequently passed to the decoder to generate the full motion sequence.
Training Objective.
To guide the hierarchical quantization process toward informative and stable motion representations, we jointly optimize a multi-term objective designed to promote accurate reconstruction, effective codebook usage, and numerical stability. The overall loss function is defined as:
where denotes the stop-gradient operator. The first term encourages accurate motion reconstruction, the second enforces codebook commitment, and the third stabilizes training by constraining the scale and bias parameters to remain close to identity mappings.
To preserve fine-grained temporal details throughout the encoding and decoding process, we employ stride-1 convolutions in both the encoder and decoder. This design avoids excessive downsampling, which can introduce high-frequency aliasing artifacts and degrade motion continuity.
Learnable scaling allows the quantization layers to focus on encoding residuals at different abstraction levels, with deeper layers specializing in finer motion refinements. Such a design increases the expressiveness and interpretability of the latent representation, contributing to higher reconstruction quality and better hierarchical decomposition of motion dynamics.
2.2 RQHC-Transformer: Residual Quantized Hierarchical Causal Transformer
multi-level token representation via MoSA-VQ, the next challenge is to generate these tokens from a textual prompt in a temporally coherent and semantically aligned manner. To this end, we propose the RQHC-Transformer, a specialized autoregressive model designed to synthesize motion token sequences across multiple quantization levels in a single unified framework.
Unlike prior transformer-based models that process motion tokens at a single resolution or require separate modules per layer [2], our RQHC-Transformer embraces the hierarchical structure of the quantized latent space. It sequentially predicts tokens from coarse to fine levels, enabling the model to first capture the overall motion structure and then refine it with increasing detail. This mirrors the residual nature of MoSA-VQ and preserves the interpretability of each quantization stage.
Input Construction.
To generate the token sequence at quantization level , we construct an input sequence that conditions on both the textual prompt and the residual context from lower levels:
where is the CLIP-based text prompt embedding, and is the embedding representing the current quantization level. The sequence is computed by aggregating token embeddings across all previous levels at each position:
Here, denotes the number of motion tokens (i.e., temporal steps) in the sequence. Each is the -th quantized token at layer , and maps each token to its corresponding embedding vector.
Relative Positional Encodings. To effectively handle long motion sequences, we incorporate a relative positional encoding scheme into our causal attention layers. Compared to absolute positional encodings, this approach better preserves attention consistency for varying sequence lengths and enhances the model’s ability to capture long-range dependencies.
Given an input token sequence, the attention score between token and token is computed as:
where and are the query and key embeddings at position and , and is a learned relative position embedding, and and are learnable global bias vectors.
This formulation allows the model to incorporate relative distance information into the attention computation. The final output of the masked self-attention is:
where is the attention score matrix incorporating relative positional encoding at layer of the -th quantization level. denotes the value matrix at the same layer.
Autoregressive Objective.
We use a multi-layer conditional maximum likelihood objective:
Here, is the batch size, is the sequence length for the -th sample, and is the number of quantization levels. denotes the ground truth discrete token at position in layer , while represents the model’s predicted context (e.g., previously generated tokens or hidden states) up to position for sample and layer . The conditional input encodes auxiliary information such as the text prompt, and denotes the model’s predicted probability distribution parameterized by weights .
2.3 Text Condition Alignment for Improved Motion Decoding
Despite the RQHC-Transformer being trained on structured datasets such as HumanML3D, real-world user prompts at inference time often exhibit substantial variation in phrasing, granularity, and semantic clarity. This distribution shift between training and deployment introduces a critical challenge: the model may underperform on out-of-distribution inputs due to linguistic mismatch between test-time instructions and training-time prompts.
To mitigate this discrepancy, we propose TCA (text condition alignment), a lightweight inference-time module that reformulates raw user instructions into a training-compatible form without altering the model’s weights. TCA operates by leveraging a pretrained language model , which serves as a prompt rewriting transformer to align free-form inputs with the motion model’s learned linguistic prior.
We denote the space of unstructured user prompts as , which includes arbitrary, potentially noisy natural language descriptions. In contrast, denotes the domain of structured, semantically aligned prompts used during training. Given a test-time input , TCA produces an aligned instruction by querying the rewriting model under a style prior :
where encapsulates syntactic and semantic formatting patterns derived from the training corpus.
The aligned prompt is then directly used as the conditioning input for the motion generation pipeline, replacing the original . This inference-time alignment improves compatibility between the textual condition and the model’s learned text-motion mapping, enabling more robust, coherent, and semantically accurate generation—particularly in zero- and few-shot settings. Since TCA operates independently of model parameters and requires no additional supervision, it offers a practical and efficient solution for open-domain deployment.
3 Experiments
We comprehensively evaluate our MOGO framework on text-to-motion generation tasks, focusing on motion quality, text-motion alignment, and generalization. Additional experimental results and ablation studies are provided in the appendix for further analysis.
3.1 Datasets and Evaluation Metrics
Datasets. We train and evaluate MOGO on HumanML3D [13] and KIT-ML [14]. HumanML3D contains 14,616 motion clips and 44,970 textual annotations, while KIT-ML includes 3,911 motion clips and 6,278 descriptions. Data splits follow T2M [13] with ratios of 0.8/0.15/0.05. To assess out-of-distribution generalization, we evaluate zero-shot performance on the CMP dataset [15], which includes 8,700 non-daily, combat-style motions. CMP provides three levels of text descriptions, and we use its test split. During CMP evaluation, prompt engineering is disabled to ensure fairness.
Evaluation Metrics. We adopt standard metrics from [13]. FID (Frechet Inception Distance) measures motion realism and serves as a primary evaluation metric. R-Precision evaluates text-motion alignment (Top-1, 2, 3). MM-Dist computes the distance between motion and text embeddings. MultiModality measures the variance of motions generated from the same text prompt.
3.2 Real-time Inference
To enhance zero- and few-shot generalization, we leverage GLM-4 [16] to rewrite user prompts that deviate from the HumanML3D style, aligning them more closely with the model’s learned text-motion priors. Moreover, benefiting from MOGO’s autoregressive design, motion tokens are generated in a streaming fashion—each token is predicted and emitted sequentially—enabling low-latency, frame-by-frame output without the need to wait for the entire sequence.
This architectural advantage allows MOGO to operate in real-time settings, achieving an inference speed of 30 frames per second on a single NVIDIA A100 GPU, making it well-suited for interactive applications such as virtual avatars, animation systems, and embodied agents.
3.3 Comparison to State-of-the-art Approaches
3.3.1 Reconstruction Quality of Motion Encoder
We compare the reconstruction fidelity of our MoSA-VQ with other state-of-the-art motion VAEs. As shown in Table 1, our encoder achieves the lowest FID on both HumanML3D and KIT-ML datasets.
| HumanML3D | KIT-ML | ||
| Method | FID | Method | FID |
| M2DM [17] | 0.063±0.001 | M2DM [17] | 0.413±0.009 |
| T2M-GPT [3] | 0.070±0.001 | T2M-GPT [3] | 0.472±0.011 |
| MoMask [2] | 0.019±0.001 | MoMask [2] | 0.112±0.002 |
| MMM [4] | 0.075±0.001 | MMM [4] | 0.641±0.014 |
| MOGO | 0.013±0.001 | MOGO | 0.037±0.001 |
3.3.2 Quantitative Results
As shown in Table 2, MOGO exhibits consistently strong performance across the HumanML3D, KIT-ML, and CMP (zero-shot) benchmarks. With the incorporation of our proposed TCA module, MOGO achieves leading results in several key metrics—particularly in terms of generation quality and diversity—while maintaining competitive performance across all others.
On in-distribution datasets like HumanML3D and KIT-ML, MOGO with TCA demonstrates top-tier performance across most metrics, though it exhibits a slight increase in FID compared to the base MOGO, the overall generation quality remains state-of-the-art, underscoring the robustness of our hierarchical architecture.
| Datasets | Methods | R Precision | FID | MultiModal Dist | MultiModality | ||
| Top 1 | Top 2 | Top 3 | |||||
| HumanML3D | MotionDiffuse [18] | 0.4910.001 | 0.6810.001 | 0.7820.001 | 0.6300.001 | 3.1130.001 | 1.5530.042 |
| T2M-GPT† [3] | 0.4910.003 | 0.6800.002 | 0.7750.002 | 0.1160.004 | 3.1180.011 | 1.8560.011 | |
| Fg-T2M [19] | 0.4920.002 | 0.6830.003 | 0.7830.003 | 0.2430.019 | 3.1090.007 | 1.6140.049 | |
| AttT2M† [5] | 0.4990.005 | 0.6900.006 | 0.7860.004 | 0.1120.004 | 3.0380.016 | 2.4520.043 | |
| MotionGPT† [20] | 0.4920.003 | 0.6810.003 | 0.7780.002 | 0.2320.008 | 3.0960.009 | 2.0080.084 | |
| MoMask [2] | 0.5210.002 | 0.7130.002 | 0.8070.002 | 0.0450.002 | 2.9580.008 | 1.2410.040 | |
| MMM [4] | 0.5040.002 | 0.696 0.003 | 0.7940.004 | 0.0800.004 | 2.9980.007 | 1.2260.035 | |
| MOGO† | 0.5150.003 | 0.7090.003 | 0.8010.003 | 0.0380.002 | 2.9510.008 | 2.1080.070 | |
| MOGO with TCA† | 0.5270.007 | 0.7220.008 | 0.8270.012 | 0.0640.003 | 2.8490.003 | 2.3440.037 | |
| KIT-ML | MotionDiffuse [18] | 0.4170.004 | 0.6210.004 | 0.7390.004 | 1.9540.062 | 2.9580.005 | 0.7300.013 |
| T2M-GPT† [3] | 0.4160.006 | 0.6270.006 | 0.7450.006 | 0.5140.029 | 3.0070.023 | 1.5700.039 | |
| Fg-T2M [19] | 0.4180.005 | 0.6260.004 | 0.7450.004 | 0.5710.047 | 3.1140.015 | 1.0190.029 | |
| AttT2M† [5] | 0.4130.006 | 0.6320.006 | 0.7510.006 | 0.8700.039 | 3.0390.016 | 2.2810.043 | |
| MotionGPT† [20] | 0.3660.005 | 0.5580.004 | 0.6800.005 | 0.5100.004 | 3.5270.021 | 2.3280.117 | |
| MoMask [2] | 0.4330.007 | 0.6560.005 | 0.7810.005 | 0.2040.011 | 2.7790.022 | 1.1310.043 | |
| MMM [4] | 0.3810.005 | 0.5900.006 | 0.7180.005 | 0.4290.019 | 3.1460.019 | 1.1050.026 | |
| MOGO† | 0.4200.007 | 0.6340.007 | 0.7540.007 | 0.1910.016 | 2.9570.029 | 2.0630.066 | |
| MOGO with TCA† | 0.4470.023 | 0.6680.016 | 0.8010.007 | 0.3130.009 | 2.8490.007 | 2.2730.073 | |
| CMP (zero-shot) | T2M-GPT† [3] | 0.0610.003 | 0.1030.005 | 0.1470.006 | 16.0920.099 | 4.1790.049 | 2.1180.033 |
| AttT2M† [5] | 0.0650.004 | 0.1090.008 | 0.1470.008 | 18.4030.071 | 4.0480.017 | 2.2080.019 | |
| MotionGPT† [20] | 0.0500.002 | 0.0940.002 | 0.1330.003 | 15.6540.183 | 4.4310.021 | 5.5350.259 | |
| MoMask [2] | 0.0620.003 | 0.1080.005 | 0.1500.004 | 24.3510.205 | 4.8170.022 | 1.6510.050 | |
| MMM [4] | 0.0670.004 | 0.1160.008 | 0.1540.008 | 17.0870.313 | 4.3600.017 | 2.8020.011 | |
| MOGO† | 0.0710.003 | 0.1240.004 | 0.1830.004 | 10.3880.171 | 3.8470.029 | 4.5620.066 | |
| MOGO with TCA† | 0.1220.006 | 0.2270.011 | 0.3040.004 | 6.8730.073 | 3.0400.014 | 4.2990.013 | |
Red = Best, Blue = Second Best, orange = Third Best.
This advantage becomes more pronounced in the CMP zero-shot setting, where other transformer-based approaches often suffer from significant degradation. In contrast, MOGO with TCA achieves the best results in multiple metrics, highlighting its ability to generalize effectively to out-of-distribution prompts. These improvements stem from the synergy between our hierarchical generation framework and the inference-time input alignment mechanism introduced by TCA.
Furthermore, MOGO supports streaming token-wise generation, allowing for real-time inference without sacrificing output quality. This combination of strong generalization, high overall performance, and low-latency generation makes MOGO well-suited for practical deployment in open-ended, real-world motion generation scenarios.
3.3.3 Qualitative Comparison and Long-range Generation
Qualitative Comparison. As shown in Figure 3, our generated motions are smoother, more consistent and better aligned with input prompts compared to other methods (more results are shown in Appendix A, and videos are also provided in the supplementary material). Built upon a hierarchical architecture, MOGO enables longer motion generation without sacrificing coherence or realism, demonstrating competitive performance among transformer-based autoregressive models and strong generalization capabilities (more ablation studies are shown in Appendix).
4 Related Work
Human Motion Generation. Recent advances in human motion generation have enabled conditioning on modalities like text, audio, music, and images [1]. Early deterministic models [21, 22] often produced over-smoothed, unrealistic motions. To address this, stochastic approaches, such as GANs [23, 24] and VAE-based models [25, 26], improved motion diversity. Text-to-motion generation gained prominence with works like [13], which used temporal VAEs to model text-motion distributions. Recently, diffusion-based methods [27, 18, 28, 17, 8, 6] and transformer-based approaches [2, 3, 20] have led the field.
LLM-Based Motion Generation Models. LLM-based architectures have become a cornerstone of text-to-motion generation, leveraging their ability to model sequential data and adapt to varied tasks [2, 3, 4, 5, 20]. Models like MoMask [2] and MMM [4] use masked token modeling to produce high-quality motions but struggle with streaming output and generalization in low-data or out-of-distribution settings due to their bidirectional design. Conversely, autoregressive models like T2M-GPT [3] and AttT2M [5] enable sequential generation, making them suitable for real-time applications and scalable with larger datasets, though they often sacrifice some motion quality. Efforts like MotionGPT [20] integrate multimodal language modeling but face challenges in achieving high-fidelity motion outputs. Our MOGO framework addresses these issues by combining efficient encoding through the MoSA-VQ with a single-pass autoregressive transformer, ensuring both high-quality motion and streaming capabilities.
Hierarchical Transformers. Hierarchical transformer architectures excel in domains like NLP [12, 29], image generation [30], and vision tasks [31, 32]. By processing data at multiple abstraction levels, these models enhance representation capacity and scalability. For instance, Swin Transformers [31] support scalable high-resolution vision, while CogView2 [30] enables high-fidelity image synthesis. In motion generation, hierarchical transformers are underexplored, especially for autoregressive frameworks with residual quantization. Our RQHC-Transformer in MOGO leverages hierarchical modeling to efficiently process multi-layer motion tokens, improving quality and generalization.
Motion Tokenization. Discretizing continuous motion data into tokens v‘ia vector quantization is central to transformer-based motion generation. TM2T [33] introduced VQ-VAE to map motions to discrete sequences. T2M-GPT [3] enhanced token quality with exponential moving average and codebook reset techniques. AttT2M [5] improved quantization through body-part-aware encoding. MoMask [2] advanced this with residual vector quantization (RVQ), producing multi-level tokens for better reconstruction quality. Compared to prior work, our approach introduces learnable feature scaling into the residual vector quantization process, allowing each quantization level to adaptively adjust the magnitude of its residuals. This scaled RVQ mechanism stabilizes the residual distribution across quantization stages, mitigates the risk of later stages collapsing to noise, and ensures more effective codebook utilization. As a result, the model benefits from improved reconstruction fidelity, better token expressiveness, and enhanced training stability.
5 Limitations and Discussion
Despite MOGO’s strong performance in real-time streaming and zero-shot generalization, several limitations remain.
Motion Editing. Due to its unidirectional autoregressive design, MOGO struggles with temporal editing tasks such as infilling or interpolation. This limits its use in scenarios like animation post-processing, where flexible motion manipulation is needed.
Generation Length. While relative positional encoding allows MOGO to generate up to 260 frames in cyclic motions, non-continuous sequences remain limited to around 196 frames. This is primarily due to weak temporal correlations in motion data, which hinder the use of segment-level memory mechanisms like Transformer-XL [34].
To address these issues, future work should focus on (1) enabling conditional adjustments after initial generation to support dynamic prompts, and (2) improving local condition alignment for seamless mid-sequence control. These enhancements could make MOGO more adaptable for interactive and editable motion generation.
6 Conclusion
We presented MOGO, a one-pass autoregressive framework for high-quality and real-time 3D human motion generation. By combining the MoSA-VQ module for scale-adaptive residual vector quantization with the RQHC-Transformer for hierarchical causal decoding, MOGO achieves efficient and temporally coherent motion synthesis. The integration of TCA mechanism further enhances semantic alignment and enables better generalization in zero- and few-shot scenarios. Our framework supports low-latency, frame-by-frame inference, making it well-suited for real-time and streamable applications. Extensive experiments demonstrate that MOGO outperforms prior transformer-based methods across multiple benchmarks in terms of generation quality and text-motion alignment. While current limitations remain in motion editing and handling complex non-cyclic sequences, MOGO provides a robust foundation for future research toward controllable, editable, and interactive motion generation.
Appendix A Generation Results
Stylized Motion Generation.
Figure 4 illustrates the results of our model applied to stylized motion generation. Three representative action types are shown, each demonstrating the model’s ability to preserve motion semantics while adapting to distinct style patterns. To further showcase the generality and diversity of our approach, we provide additional stylized motion sequences in MP4 format in the supplementary material.
Qualitative Comparison.
Figures 6–9 present a qualitative comparison of motion generation results given the same text prompt, showcasing outputs among our model, Momask [2] and MotionGPT [20]. As observed, MOGO produces more complete and coherent motion sequences, particularly excelling in motion completeness and fidelity. For a more intuitive evaluation, we additionally provide corresponding video results in the supplementary material (MP4 format), which further highlight the differences in motion quality and realism.
Appendix B Ablation Study
Impact of Learnable Scaling in MoSA-VQ.
We investigate the role of the proposed learnable scaling mechanism in MoSA-VQ by comparing models with and without this design. In the baseline configuration, residual features at each quantization level are directly quantized without any adaptive normalization. In contrast, our full model applies a layer-specific affine transformation before quantization, allowing each level to adjust the residual space dynamically.
As shown in Table 3, incorporating learnable scale and bias parameters leads to consistent improvements across multiple evaluation metrics. The model demonstrates better reconstruction quality, enhanced alignment between motion and text, and stronger semantic consistency. These results highlight the benefit of introducing adaptive normalization in residual quantization, which facilitates more expressive and stable token representations.
| Codebook size | w/ Learnable Scaling | ||
| FID | Top1 | MM-Dist | |
| 2048 512 | 0.0520.005 | 0.4980.002 | 2.9970.008 |
| 4096 256 | 0.0470.005 | 0.5010.002 | 2.9010.008 |
| 8192 128 | 0.0380.003 | 0.5150.007 | 2.9510.003 |
| Codebook size | w/o Learnable Scaling | ||
| FID | Top1 | MM-Dist | |
| 2048 512 | 0.0940.011 | 0.4840.002 | 3.2730.009 |
| 4096 256 | 0.0910.004 | 0.4930.004 | 3.4510.017 |
| 8192 128 | 0.0790.002 | 0.5020.002 | 3.0020.008 |
Effect of Quantization Depth in MoSA-VQ.
We further explore how the number of residual quantization layers in MoSA-VQ affects the model’s overall performance. As residual quantization hierarchically captures motion information from coarse to fine granularity, increasing the depth allows more detailed refinement of the motion representation.
| Layers | Reconstruction | Generation | ||||
| FID | Top1 | MM-Dist | FID | Top1 | MM-Dist | |
| 1 | 0.0670.001 | 0.5080.002 | 2.9970.007 | 0.0680.009 | 0.4690.002 | 3.1380.006 |
| 2 | 0.0460.001 | 0.5040.001 | 2.9840.010 | 0.0530.008 | 0.4670.003 | 3.1140.002 |
| 3 | 0.0350.001 | 0.5080.003 | 2.9800.004 | 0.0440.005 | 0.4980.002 | 2.9970.008 |
| 4 | 0.0230.001 | 0.5050.003 | 2.9890.006 | 0.0450.005 | 0.5010.002 | 2.9820.002 |
| 5 | 0.0170.001 | 0.5110.001 | 2.9800.003 | 0.0410.005 | 0.4980.002 | 2.9770.002 |
| 6 | 0.0130.001 | 0.5100.001 | 2.9820.006 | 0.0380.003 | 0.5150.002 | 2.9510.003 |
| 7 | 0.0130.001 | 0.5030.002 | 2.9930.007 | 0.0400.003 | 0.5130.007 | 2.9430.003 |
As shown in Table 4, models with a shallow quantization depth tend to underperform, suggesting that limited hierarchy constrains the representational capacity of the latent code. As the number of layers increases, performance improves consistently across reconstruction fidelity, text-motion alignment, and multimodal distance. However, the gains tend to plateau beyond a certain depth, indicating diminishing returns. This analysis supports the use of a moderate number of residual layers as a trade-off between expressiveness and efficiency.
Codebook Size. We evaluate various codebook sizes on HumanML3D. As shown in Table 5, achieves the best FID for generation and competitive reconstruction.
Dataset Size. We evaluate MOGO’s scalability by mixing HumanML3D with varying proportions of CMP data (Table 6). Without retraining the MoSA-VQ, generation quality improves significantly with increased data volume.
| Codebook Size | Reconstruction | Generation | ||||
| FID | Top1 | MM-Dist | FID | Top1 | MM-Dist | |
| 512 512 | 0.0220.001 | 0.5080.003 | 2.9970.007 | 0.2030.009 | 0.4690.002 | 3.1380.006 |
| 1024 1024 | 0.0150.001 | 0.5110.002 | 2.9840.010 | 0.1140.008 | 0.4670.003 | 3.1140.009 |
| 2048 512 | 0.0170.001 | 0.5110.003 | 2.9800.007 | 0.0520.005 | 0.4980.002 | 2.9970.008 |
| 4096 256 | 0.0190.001 | 0.5100.002 | 2.9890.006 | 0.0470.005 | 0.5010.002 | 2.9010.008 |
| 8192 128 | 0.0130.001 | 0.5100.002 | 2.9890.007 | 0.0380.003 | 0.5150.007 | 2.9510.003 |
| Epochs | w/ 0% CMP | w/ 50% CMP | w/ 100% CMP |
| 50 | 0.671 | 0.601 | 0.501 |
| 80 | 0.437 | 0.379 | 0.284 |
| 100 | 0.347 | 0.263 | 0.205 |
Input Condition. We compared the performance of input conditions by either adding the prompt token and layer token as a sequence prefix or not. As shown in Table 7, our experiments reveal that when without PnQ (adding prompt token to quantization layer token) is applied, the correlation between generated motion sequences and text (measured by R-Precision and MM Distance) decreases significantly. This may result from an additional layer token between action sequence tokens and prompt tokens during training, leading to some attention loss, as shown in Figure 10.
| Codebook size | w/ PnQ | ||
| FID | Top1 | MM-Dist | |
| 2048 512 | 0.0520.005 | 0.4980.002 | 2.9970.008 |
| 4096 256 | 0.0470.005 | 0.5010.002 | 2.9010.008 |
| 8192 128 | 0.0380.003 | 0.5150.007 | 2.9510.003 |
| Codebook size | w/o PnQ | ||
| FID | Top1 | MM-Dist | |
| 2048 512 | 0.0960.005 | 0.4520.002 | 3.2930.009 |
| 4096 256 | 0.0910.005 | 0.4590.002 | 3.1410.007 |
| 8192 128 | 0.0850.002 | 0.4820.003 | 3.1270.008 |
Appendix C Algorithm
The core training code structure of our MoSA-VQ model is illustrated below. The variable abbreviations used in the code are defined as follows:
The core training code structure of our RQHC-Tranformer is illustrated below. The variable abbreviations used in the code are defined as follows:

### v2

@@
-MOGO: Residual Quantized Hierarchical Causal Transformer for High-Quality and Real-Time 3D Human Motion Generation
+MOGO: Residual Quantized Hierarchical Causal Transformer for Real-Time and Infinite-Length 3D Human Motion Generation
 Abstract
-Recent advances in transformer-based text-to-motion generation have led to impressive progress in synthesizing high-quality human motion.
-Nevertheless, jointly achieving high fidelity, streaming capability, real-time responsiveness, and scalability remains a fundamental challenge.
-In this paper, we propose MOGO (Motion Generation with One-pass), a novel autoregressive framework tailored for efficient and real-time 3D motion generation.
-MOGO comprises two key components: (1) MoSA-VQ, a motion scale-adaptive residual vector quantization module that hierarchically discretizes motion sequences with learnable scaling to produce compact yet expressive representations; and (2) RQHC-Transformer, a residual quantized hierarchical causal transformer that generates multi-layer motion tokens in a single forward pass, significantly reducing inference latency.
-To enhance semantic fidelity, we further introduce a text condition alignment mechanism that improves motion decoding under textual control.
-Extensive experiments on benchmark datasets including HumanML3D, KIT-ML, and CMP demonstrate that MOGO achieves competitive or superior generation quality compared to state-of-the-art transformer-based methods, while offering substantial improvements in real-time performance, streaming generation, and generalization under zero-shot settings.
-1 Introduction
+Recent advances in transformer-based text-to-motion generation have significantly improved motion quality.
+However, achieving both real-time performance and long-horizon scalability remains an open challenge.
+In this paper, we present MOGO (Motion Generation with One-pass), a novel autoregressive framework for efficient and scalable 3D human motion generation.
+MOGO consists of two key components.
+First, we introduce MoSA-VQ, a motion scale-adaptive residual vector quantization module that hierarchically discretizes motion sequences through learnable scaling parameters, enabling dynamic allocation of representation capacity and producing compact yet expressive multi-level representations.
+Second, we design the RQHC-Transformer, a residual quantized hierarchical causal transformer that decodes motion tokens in a single forward pass.
+Each transformer block aligns with one quantization level, allowing hierarchical abstraction and temporally coherent generation with strong semantic flow.
+Compared to diffusion- and LLM-based approaches, MOGO achieves lower inference latency while preserving high motion fidelity.
+Moreover, its hierarchical latent design enables seamless and controllable infinite-length motion generation, with stable transitions and the ability to adaptively incorporate updated control signals at arbitrary points in time.
+To further enhance generalization and interpretability, we introduce Textual Condition Alignment (TCA), which leverages large language models with Chain-of-Thought reasoning to bridge the gap between real-world prompts and training data.
+TCA not only improves zero-shot performance on unseen datasets but also enriches motion comprehension for in-distribution prompts through explicit intent decomposition.
+Extensive experiments on HumanML3D, KIT-ML, and the unseen CMP dataset demonstrate that MOGO outperforms prior methods in generation quality, inference efficiency, and temporal scalability.
+I Introduction
 Text-to-motion generation is a rapidly evolving research area with growing importance in virtual environments, such as gaming, AR/VR, and humanoid robotics [1].
@@
 Diffusion models often rely on iterative refinement processes, which introduce significant inference latency and hinder their suitability for real-time or interactive applications [7].
-On the other hand, LLM-based models, although autoregressive, typically involve large parameter sizes and long context dependencies, leading to high memory and computation costs that challenge deployment on lightweight or streaming scenarios [9].
-A broader discussion of related research trends is provided in Section 4.
-To address these issues, we propose MOGO (Motion Generation with One-pass), a transformer-based autoregressive framework that directly predicts motion tokens in a single forward pass.
-Compared to diffusion-based models, MOGO enables streamable motion decoding without iterative sampling, significantly reducing inference time.
-Furthermore, unlike LLM-based motion generators, MOGO adopts a lightweight yet expressive transformer architecture with considerably fewer parameters—allowing for real-time token generation while maintaining high motion fidelity.
-MOGO is composed of two key components: (1) MoSA-VQ (Motion Scale-Adaptive Residual VQ-VAE), which discretizes motion sequences via hierarchical quantization and learnable scaling, producing compact and adaptive representations [11]; and (2) RQHC-Transformer (Residual Quantized Hierarchical Causal Transformer), which autoregressively generates multi-layer motion tokens in a single pass [12].
-As shown in Figure 1, this design allows for the synthesis of coherent and extended motion sequences while preserving streaming capability and generalization across diverse prompts and datasets.
-Extensive experiments demonstrate that MOGO achieves superior results on both in-distribution and out-of-distribution scenarios, outperforming state-of-the-art LLM-based motion generators in quality, inference speed, and scalability—making it a robust and real-time-capable solution for practical text-to-motion generation.
-Our main contributions:
-MOGO: A unified transformer-based framework for one-pass high-fidelity motion generation.
-It achieves generation quality comparable to state-of-the-art LLM- and diffusion-based methods, while significantly improving real-time efficiency and enabling streamable motion synthesis.
-MoSA-VQ: A motion discretization module that employs hierarchical quantization and learnable feature scaling for compact and expressive representations.
-RQHC-Transformer: A lightweight hierarchical causal transformer that autoregressively generates multi-layer motion tokens in one forward pass.
-All code and demonstration pages will be made publicly available on our GitHub repository upon acceptance.
-2 Methods
-We introduce MOGO, a unified framework designed for efficient and expressive text-to-motion generation.
-Mogo could generate a high-quality motion sequence from a textual description, where each frame represents a -dimensional human pose.
-As shown in Figure 2, our MOGO framework consists of two core components: (1) MoSA-VQ (left) that discretizes motion into multi-level tokens with high precision, and (2) RQHC-Transformer (right) leveraging hierarchical causal forward computation to generate motion tokens collectively in a streamlined manner (algorithms are shown in Appendix C).
-To enhance zero/few-shot performance, we incorporate text condition alignment (TCA) for improved motion decoding.
-2.1 MoSA-VQ: Motion Scale-Adaptive Residual VQ-VAE
-To effectively transform continuous human motion sequences into discrete representations suitable for autoregressive modeling, we adopt a residual vector quantized variational autoencoder (RVQ-VAE) architecture.
-This module, termed MoSA-VQ, serves as the foundation for learning compact yet expressive motion codes.
-Residual quantization enables progressive refinement of motion features across multiple levels, capturing both coarse and fine-grained temporal dynamics essential for high-fidelity generation.
-However, previous RVQ-based approaches [2, 5] often rely on manually calibrated feature ranges, which limit their adaptability and quantization efficiency in complex motion settings.
-To address this, we introduce a novel learnable feature scaling mechanism, allowing the quantization space to dynamically adapt at each hierarchy level.
-This design not only eliminates the need for manual tuning but also improves reconstruction accuracy and codebook utilization.
-Hierarchical Quantization with Feature Scaling.
-Motion encoder outputs a latent representation , where is the number of motion tokens and is the feature dimension.
-At each level , we perform feature scaling before residual quantization.
-Let be the residual feature at level 0.
-We first compute a scaled version of the residual feature:
-where and are learnable scale and bias parameters. The input is the residual feature passed from the previous quantization level and represents the information that has not yet been captured.
-We then apply vector quantization to the scaled residual:
-where maps each input to the nearest codebook entry. This quantized output exists in the scaled feature space, so we invert the transformation to project it back into the original residual space:
-The recovered quantized vector captures the portion of the motion representation explained at level , and we update the residual for the next layer by subtracting this contribution:
-This process repeats hierarchically across all levels, resulting in a set of quantized vectors . The final quantized latent representation is reconstructed by summing contributions from all layers:
-which is subsequently passed to the decoder to generate the full motion sequence.
-Training Objective.
-To guide the hierarchical quantization process toward informative and stable motion representations, we jointly optimize a multi-term objective designed to promote accurate reconstruction, effective codebook usage, and numerical stability.
-The overall loss function is defined as:
-where denotes the stop-gradient operator.
-The first term encourages accurate motion reconstruction, the second enforces codebook commitment, and the third stabilizes training by constraining the scale and bias parameters to remain close to identity mappings.
-To preserve fine-grained temporal details throughout the encoding and decoding process, we employ stride-1 convolutions in both the encoder and decoder.
-This design avoids excessive downsampling, which can introduce high-frequency aliasing artifacts and degrade motion continuity.
-Learnable scaling allows the quantization layers to focus on encoding residuals at different abstraction levels, with deeper layers specializing in finer motion refinements.
-Such a design increases the expressiveness and interpretability of the latent representation, contributing to higher reconstruction quality and better hierarchical decomposition of motion dynamics.
-2.2 RQHC-Transformer: Residual Quantized Hierarchical Causal Transformer
-multi-level token representation via MoSA-VQ, the next challenge is to generate these tokens from a textual prompt in a temporally coherent and semantically aligned manner.
-To this end, we propose the RQHC-Transformer, a specialized autoregressive model designed to synthesize motion token sequences across multiple quantization levels in a single unified framework.
-Unlike prior transformer-based models that process motion tokens at a single resolution or require separate modules per layer [2], our RQHC-Transformer embraces the hierarchical structure of the quantized latent space.
-It sequentially predicts tokens from coarse to fine levels, enabling the model to first capture the overall motion structure and then refine it with increasing detail.
-This mirrors the residual nature of MoSA-VQ and preserves the interpretability of each quantization stage.
-Input Construction.
-To generate the token sequence at quantization level , we construct an input sequence that conditions on both the textual prompt and the residual context from lower levels:
-where is the CLIP-based text prompt embedding, and is the embedding representing the current quantization level.
-The sequence is computed by aggregating token embeddings across all previous levels at each position:
-Here, denotes the number of motion tokens (i.e., temporal steps) in the sequence. Each is the -th quantized token at layer , and maps each token to its corresponding embedding vector.
-Relative Positional Encodings.
-To effectively handle long motion sequences, we incorporate a relative positional encoding scheme into our causal attention layers.
-Compared to absolute positional encodings, this approach better preserves attention consistency for varying sequence lengths and enhances the model’s ability to capture long-range dependencies.
-Given an input token sequence, the attention score between token and token is computed as:
-where and are the query and key embeddings at position and , and is a learned relative position embedding, and and are learnable global bias vectors.
-This formulation allows the model to incorporate relative distance information into the attention computation. The final output of the masked self-attention is:
-where is the attention score matrix incorporating relative positional encoding at layer of the -th quantization level. denotes the value matrix at the same layer.
-Autoregressive Objective.
-We use a multi-layer conditional maximum likelihood objective:
-Here, is the batch size, is the sequence length for the -th sample, and is the number of quantization levels.
-denotes the ground truth discrete token at position in layer , while represents the model’s predicted context (e.g., previously generated tokens or hidden states) up to position for sample and layer .
-The conditional input encodes auxiliary information such as the text prompt, and denotes the model’s predicted probability distribution parameterized by weights .
-2.3 Text Condition Alignment for Improved Motion Decoding
-Despite the RQHC-Transformer being trained on structured datasets such as HumanML3D, real-world user prompts at inference time often exhibit substantial variation in phrasing, granularity, and semantic clarity.
-This distribution shift between training and deployment introduces a critical challenge: the model may underperform on out-of-distribution inputs due to linguistic mismatch between test-time instructions and training-time prompts.
-To mitigate this discrepancy, we propose TCA (text condition alignment), a lightweight inference-time module that reformulates raw user instructions into a training-compatible form without altering the model’s weights.
-TCA operates by leveraging a pretrained language model , which serves as a prompt rewriting transformer to align free-form inputs with the motion model’s learned linguistic prior.
-We denote the space of unstructured user prompts as , which includes arbitrary, potentially noisy natural language descriptions.
-In contrast, denotes the domain of structured, semantically aligned prompts used during training.
-Given a test-time input , TCA produces an aligned instruction by querying the rewriting model under a style prior :
-where encapsulates syntactic and semantic formatting patterns derived from the training corpus.
-The aligned prompt is then directly used as the conditioning input for the motion generation pipeline, replacing the original .
-This inference-time alignment improves compatibility between the textual condition and the model’s learned text-motion mapping, enabling more robust, coherent, and semantically accurate generation—particularly in zero- and few-shot settings.
-Since TCA operates independently of model parameters and requires no additional supervision, it offers a practical and efficient solution for open-domain deployment.
-3 Experiments
-We comprehensively evaluate our MOGO framework on text-to-motion generation tasks, focusing on motion quality, text-motion alignment, and generalization.
-Additional experimental results and ablation studies are provided in the appendix for further analysis.
-3.1 Datasets and Evaluation Metrics
-Datasets.
-We train and evaluate MOGO on HumanML3D [13] and KIT-ML [14].
-HumanML3D contains 14,616 motion clips and 44,970 textual annotations, while KIT-ML includes 3,911 motion clips and 6,278 descriptions.
-Data splits follow T2M [13] with ratios of 0.8/0.15/0.05.
-To assess out-of-distribution generalization, we evaluate zero-shot performance on the CMP dataset [15], which includes 8,700 non-daily, combat-style motions.
-CMP provides three levels of text descriptions, and we use its test split.
-During CMP evaluation, prompt engineering is disabled to ensure fairness.
-Evaluation Metrics.
-We adopt standard metrics from [13].
-FID (Frechet Inception Distance) measures motion realism and serves as a primary evaluation metric.
-R-Precision evaluates text-motion alignment (Top-1, 2, 3).
-MM-Dist computes the distance between motion and text embeddings.
-MultiModality measures the variance of motions generated from the same text prompt.
-3.2 Real-time Inference
-To enhance zero- and few-shot generalization, we leverage GLM-4 [16] to rewrite user prompts that deviate from the HumanML3D style, aligning them more closely with the model’s learned text-motion priors.
-Moreover, benefiting from MOGO’s autoregressive design, motion tokens are generated in a streaming fashion—each token is predicted and emitted sequentially—enabling low-latency, frame-by-frame output without the need to wait for the entire sequence.
-This architectural advantage allows MOGO to operate in real-time settings, achieving an inference speed of 30 frames per second on a single NVIDIA A100 GPU, making it well-suited for interactive applications such as virtual avatars, animation systems, and embodied agents.
-3.3 Comparison to State-of-the-art Approaches
-3.3.1 Reconstruction Quality of Motion Encoder
-We compare the reconstruction fidelity of our MoSA-VQ with other state-of-the-art motion VAEs. As shown in Table 1, our encoder achieves the lowest FID on both HumanML3D and KIT-ML datasets.
-| HumanML3D | KIT-ML | ||
-| Method | FID | Method | FID |
-| M2DM [17] | 0.063±0.001 | M2DM [17] | 0.413±0.009 |
-| T2M-GPT [3] | 0.070±0.001 | T2M-GPT [3] | 0.472±0.011 |
-| MoMask [2] | 0.019±0.001 | MoMask [2] | 0.112±0.002 |
-| MMM [4] | 0.075±0.001 | MMM [4] | 0.641±0.014 |
-| MOGO | 0.013±0.001 | MOGO | 0.037±0.001 |
-3.3.2 Quantitative Results
-As shown in Table 2, MOGO exhibits consistently strong performance across the HumanML3D, KIT-ML, and CMP (zero-shot) benchmarks.
-With the incorporation of our proposed TCA module, MOGO achieves leading results in several key metrics—particularly in terms of generation quality and diversity—while maintaining competitive performance across all others.
-On in-distribution datasets like HumanML3D and KIT-ML, MOGO with TCA demonstrates top-tier performance across most metrics, though it exhibits a slight increase in FID compared to the base MOGO, the overall generation quality remains state-of-the-art, underscoring the robustness of our hierarchical architecture.
-| Datasets | Methods | R Precision | FID | MultiModal Dist | MultiModality | ||
-| Top 1 | Top 2 | Top 3 | |||||
-| HumanML3D | MotionDiffuse [18] | 0.4910.001 | 0.6810.001 | 0.7820.001 | 0.6300.001 | 3.1130.001 | 1.5530.042 |
-| T2M-GPT† [3] | 0.4910.003 | 0.6800.002 | 0.7750.002 | 0.1160.004 | 3.1180.011 | 1.8560.011 | |
-| Fg-T2M [19] | 0.4920.002 | 0.6830.003 | 0.7830.003 | 0.2430.019 | 3.1090.007 | 1.6140.049 | |
-| AttT2M† [5] | 0.4990.005 | 0.6900.006 | 0.7860.004 | 0.1120.004 | 3.0380.016 | 2.4520.043 | |
-| MotionGPT† [20] | 0.4920.003 | 0.6810.003 | 0.7780.002 | 0.2320.008 | 3.0960.009 | 2.0080.084 | |
-| MoMask [2] | 0.5210.002 | 0.7130.002 | 0.8070.002 | 0.0450.002 | 2.9580.008 | 1.2410.040 | |
-| MMM [4] | 0.5040.002 | 0.696 0.003 | 0.7940.004 | 0.0800.004 | 2.9980.007 | 1.2260.035 | |
-| MOGO† | 0.5150.003 | 0.7090.003 | 0.8010.003 | 0.0380.002 | 2.9510.008 | 2.1080.070 | |
-| MOGO with TCA† | 0.5270.007 | 0.7220.008 | 0.8270.012 | 0.0640.003 | 2.8490.003 | 2.3440.037 | |
-| KIT-ML | MotionDiffuse [18] | 0.4170.004 | 0.6210.004 | 0.7390.004 | 1.9540.062 | 2.9580.005 | 0.7300.013 |
-| T2M-GPT† [3] | 0.4160.006 | 0.6270.006 | 0.7450.006 | 0.5140.029 | 3.0070.023 | 1.5700.039 | |
-| Fg-T2M [19] | 0.4180.005 | 0.6260.004 | 0.7450.004 | 0.5710.047 | 3.1140.015 | 1.0190.029 | |
-| AttT2M† [5] | 0.4130.006 | 0.6320.006 | 0.7510.006 | 0.8700.039 | 3.0390.016 | 2.2810.043 | |
-| MotionGPT† [20] | 0.3660.005 | 0.5580.004 | 0.6800.005 | 0.5100.004 | 3.5270.021 | 2.3280.117 | |
-| MoMask [2] | 0.4330.007 | 0.6560.005 | 0.7810.005 | 0.2040.011 | 2.7790.022 | 1.1310.043 | |
-| MMM [4] | 0.3810.005 | 0.5900.006 | 0.7180.005 | 0.4290.019 | 3.1460.019 | 1.1050.026 | |
-| MOGO† | 0.4200.007 | 0.6340.007 | 0.7540.007 | 0.1910.016 | 2.9570.029 | 2.0630.066 | |
-| MOGO with TCA† | 0.4470.023 | 0.6680.016 | 0.8010.007 | 0.3130.009 | 2.8490.007 | 2.2730.073 | |
-| CMP (zero-shot) | T2M-GPT† [3] | 0.0610.003 | 0.1030.005 | 0.1470.006 | 16.0920.099 | 4.1790.049 | 2.1180.033 |
-| AttT2M† [5] | 0.0650.004 | 0.1090.008 | 0.1470.008 | 18.4030.071 | 4.0480.017 | 2.2080.019 | |
-| MotionGPT† [20] | 0.0500.002 | 0.0940.002 | 0.1330.003 | 15.6540.183 | 4.4310.021 | 5.5350.259 | |
-| MoMask [2] | 0.0620.003 | 0.1080.005 | 0.1500.004 | 24.3510.205 | 4.8170.022 | 1.6510.050 | |
-| MMM [4] | 0.0670.004 | 0.1160.008 | 0.1540.008 | 17.0870.313 | 4.3600.017 | 2.8020.011 | |
-| MOGO† | 0.0710.003 | 0.1240.004 | 0.1830.004 | 10.3880.171 | 3.8470.029 | 4.5620.066 | |
-| MOGO with TCA† | 0.1220.006 | 0.2270.011 | 0.3040.004 | 6.8730.073 | 3.0400.014 | 4.2990.013 | |
-Red = Best, Blue = Second Best, orange = Third Best.
-This advantage becomes more pronounced in the CMP zero-shot setting, where other transformer-based approaches often suffer from significant degradation.
-In contrast, MOGO with TCA achieves the best results in multiple metrics, highlighting its ability to generalize effectively to out-of-distribution prompts.
-These improvements stem from the synergy between our hierarchical generation framework and the inference-time input alignment mechanism introduced by TCA.
-Furthermore, MOGO supports streaming token-wise generation, allowing for real-time inference without sacrificing output quality.
-This combination of strong generalization, high overall performance, and low-latency generation makes MOGO well-suited for practical deployment in open-ended, real-world motion generation scenarios.
-3.3.3 Qualitative Comparison and Long-range Generation
-Qualitative Comparison.
-As shown in Figure 3, our generated motions are smoother, more consistent and better aligned with input prompts compared to other methods (more results are shown in Appendix A, and videos are also provided in the supplementary material).
-Built upon a hierarchical architecture, MOGO enables longer motion generation without sacrificing coherence or realism, demonstrating competitive performance among transformer-based autoregressive models and strong generalization capabilities (more ablation studies are shown in Appendix).
-4 Related Work
+On the other hand, LLM-based models, although autoregressive, typically involve large parameter sizes and long context dependencies, leading to high memory and computation costs that challenge deployment on lightweight scenarios [9].
+A broader discussion of related research trends is provided in Section II.
+To address the challenges of motion generation quality, inference efficiency, and generalization, we propose MOGO, a unified and efficient transformer-based framework for expressive text-to-motion synthesis.
+MOGO generates high-fidelity motion sequences from textual input in a single forward pass, combining structural compactness, temporal coherence, and extendable decoding.
+The framework is composed of four key components, each targeting a core aspect of the generation pipeline:
+We propose MoSA-VQ, a motion structure-aware quantization framework with two key innovations: (i) a learnable modulation mechanism that adaptively scales residual magnitudes across quantization levels, and (ii) a novel cross-level decorrelation loss that enforces statistical orthogonality between quantized vectors and subsequent residuals.
+Together, these components enable more compact and hierarchically disentangled representations, leading to better reconstruction under constrained codebook capacity.
+RQHC-Transformer (Residual Quantized Hierarchical Causal Transformer) functions as the motion decoder.
+It autoregressively generates multi-layer motion tokens in a single forward pass, with each transformer block dedicated to decoding one level of the quantized hierarchy.
+This design ensures temporally coherent synthesis while preserving decoding efficiency and compatibility with interactive applications.
+Infinite-Length Generation is enabled by the autoregressive nature of MOGO’s architecture.
+Unlike diffusion-based methods that require fixed-length sampling, MOGO supports seamless motion extension with consistent structure and fidelity, making it suitable for real-time or open-ended generation scenarios.
+Textual Condition Alignment (TCA) leverages large language models with Chain-of-Thought reasoning to bridge the gap between real-world user prompts and the model’s training distribution.
+Beyond simple style normalization, TCA explicitly decomposes complex motion instructions into structured semantic cues, enabling more accurate motion interpretation.
+This enhances MOGO’s generalization capability across both in-distribution and zero-/few-shot scenarios, while improving control fidelity and robustness to diverse user inputs.
+II Related Work
 Human Motion Generation.
@@
 LLM-based architectures have become a cornerstone of text-to-motion generation, leveraging their ability to model sequential data and adapt to varied tasks [2, 3, 4, 5, 22].
-Models like MoMask [2] and MMM [4] use masked token modeling to produce high-quality motions but struggle with streaming output and generalization in low-data or out-of-distribution settings due to their bidirectional design.
+Models like MoMask [2] and MMM [4] use masked token modeling to produce high-quality motions but struggle with extendable output and generalization in low-data or out-of-distribution settings due to their bidirectional design.
 Conversely, autoregressive models like T2M-GPT [3] and AttT2M [5] enable sequential generation, making them suitable for real-time applications and scalable with larger datasets, though they often sacrifice some motion quality.
 Efforts like MotionGPT [22] integrate multimodal language modeling but face challenges in achieving high-fidelity motion outputs.
-Our MOGO framework addresses these issues by combining efficient encoding through the MoSA-VQ with a single-pass autoregressive transformer, ensuring both high-quality motion and streaming capabilities.
+Our MOGO framework addresses these issues by combining efficient encoding through the MoSA-VQ with a single-pass autoregressive transformer, ensuring both high-quality motion and interactive capabilities.
 Hierarchical Transformers.
@@
 As a result, the model benefits from improved reconstruction fidelity, better token expressiveness, and enhanced training stability.
-5 Limitations and Discussion
-Despite MOGO’s strong performance in real-time streaming and zero-shot generalization, several limitations remain.
-Motion Editing.
-Due to its unidirectional autoregressive design, MOGO struggles with temporal editing tasks such as infilling or interpolation.
-This limits its use in scenarios like animation post-processing, where flexible motion manipulation is needed.
-Generation Length.
-While relative positional encoding allows MOGO to generate up to 260 frames in cyclic motions, non-continuous sequences remain limited to around 196 frames.
-This is primarily due to weak temporal correlations in motion data, which hinder the use of segment-level memory mechanisms like Transformer-XL [34].
-To address these issues, future work should focus on (1) enabling conditional adjustments after initial generation to support dynamic prompts, and (2) improving local condition alignment for seamless mid-sequence control.
-These enhancements could make MOGO more adaptable for interactive and editable motion generation.
-6 Conclusion
+III Methods
+III-A Overview of the MOGO Framework
+Our proposed framework, MOGO, as showned in Figure 1, consists of four key components designed for efficient and expressive text-to-motion generation.
+First, MoSA-VQ discretizes motion sequences into multi-level latent tokens using hierarchical residual quantization with adaptive scaling.
+Second, RQHC-Transformer autoregressively decodes the quantized representations with a stack of level-aligned transformer blocks.
+Third, MOGO supports infinite-length generation via seamless autoregressive extension.
+Lastly, we introduce TCA to enhance generalization to real-world prompts.
+We elaborate on each component in the following subsections.
+III-B MoSA-VQ: Motion Scale-Adaptive Residual VQ-VAE
+To transform continuous human motion into discrete representations suitable for autoregressive generation, we build upon residual vector quantized variational autoencoders (RVQ-VAEs) and propose MoSA-VQ.
+While standard RVQ-based approaches [2, 5] hierarchically encode motion features through fixed-stage quantization, they often rely on manually tuned feature ranges and static scaling behavior, which limit their adaptability and quantization efficiency—especially in the presence of diverse and complex motions.
+Adaptive Hierarchical Quantization.
+To address these limitations, we introduce a learnable feature scaling mechanism into each quantization level.
+Specifically, each level is equipped with its own learnable scale and bias parameters, which dynamically adjust the amplitude and offset of the input residual features before quantization.
+This adaptive design allows each level to allocate representational capacity where needed, resulting in more effective codebook usage and higher reconstruction accuracy without the need for handcrafted calibration.
+At each hierarchical quantization level , the residual vector is first transformed using a learnable affine mapping. We denote this process as:
+where is a diagonal scaling matrix, is a learnable bias vector, denotes nearest-neighbor quantization via a codebook, and represents the decoded contribution at level .
+This hierarchical quantization scheme enables progressive residual refinement.
+Unlike conventional residual vector quantization (RVQ) frameworks that apply uniform scaling across all levels, our design introduces level-specific adaptive modulation through learnable parameters and .
+This flexibility allows coarse levels to capture global structure, while deeper layers focus on subtle motion or spatial details.
+The inverse affine transformation ensures consistent feature recovery, and stride-1 convolutions in both encoder and decoder maintain temporal continuity without aliasing artifacts.
+We have
+Training Objective.
+Our full training objective combines four terms: motion reconstruction, codebook commitment, lightweight modulation regularization, and a novel cross-level decorrelation loss.
+We adopt an reconstruction loss to ensure accurate motion recovery, and a standard commitment loss to encourage the encoder residuals to remain close to their quantized counterparts .
+A mild regularization is also applied to the per-layer scaling and bias parameters , following prior work on stable quantization with learnable modulation.
+Additionally, we introduce a cross-level decorrelation loss, which explicitly penalizes statistical redundancy between the quantized vector and the subsequent residual :
+This loss is motivated by the principle that, in an ideal residual quantization hierarchy, each level should encode novel information orthogonal to what has already been captured.
+A nonzero covariance between and indicates overlapping representation subspaces, leading to inefficient coding and excessive reliance on deeper quantizers.
+By minimizing their batch-wise covariance, we enforce an approximate statistical orthogonality between levels, which encourages disentangled information flow and hierarchical efficiency.
+From an information-theoretic perspective, this decorrelation loss can be viewed as promoting minimal mutual information between consecutive encoding layers, thus maximizing the effective bit utilization of each quantized token.
+Empirically, we find this regularization yields more compact latent codes and improves reconstruction quality under constrained codebook budgets.
+The final loss function is:
+Here, , and are weight factors.
+III-C RQHC-Transformer: Residual Quantized Hierarchical Causal Transformer
+To fully leverage the high-quality hierarchical representations produced by MoSA-VQ, we propose the Residual Quantized Hierarchical Causal Transformer (RQHC-Transformer), an autoregressive decoder structurally aligned with the residual quantization process.
+In contrast to single-resolution or per-layer decoding approaches, RQHC jointly models all quantization levels in a residual-aware and level-aware manner.
+Moreover, we adopt a causal Transformer design, which—compared to bidirectional transformers, LLM-style decoders, or iterative diffusion models—offers efficient one-pass inference, supports streamable generation, and naturally enables long-horizon synthesis.
+The coarse-to-fine decoding progression preserves structural semantics, mitigates token interference, and promotes stable, coherent motion generation with improved interpretability.
+This hierarchical causal architecture progressively refines coarse motion structures, preserving semantic consistency across abstraction levels and reducing token-level interference.
+Crucially, its design is structurally aligned with the residual quantization process of MoSA-VQ, enabling stable information flow across layers.
+Together, this synergy allows MOGO to extend motion from any given frame with temporal continuity and semantic fidelity, laying the foundation for real-time, controllable infinite-length generation.
+Hierarchical Causal Generation.
+To generate the token sequence at quantization level , we construct an input sequence that conditions on both the textual prompt and the residual context from lower levels:
+where is the CLIP-based text prompt embedding, and is the embedding representing the current quantization level.
+The sequence is computed by aggregating token embeddings across all previous levels at each position:
+Here, denotes the number of motion tokens (i.e., temporal steps) in the sequence. Each is the -th quantized token at layer , and maps each token to its corresponding embedding vector.
+Relative Positional Encodings.
+To effectively handle long motion sequences, we incorporate a relative positional encoding scheme into our causal attention layers.
+Compared to absolute positional encodings, this approach better preserves attention consistency for varying sequence lengths and enhances the model’s ability to capture long-range dependencies.
+Given an input token sequence, the attention score between token and token is computed as:
+where and are the query and key embeddings at position and , and is a learned relative position embedding, and and are learnable global bias vectors. The final output of the self-attention is:
+where is the attention score matrix incorporating relative positional encoding at layer of the -th quantization level. denotes the value matrix at the same layer.
+Training Objective.
+To learn temporally coherent and hierarchically consistent token generation, we adopt a multi-layer autoregressive training objective defined over all quantization levels.
+The loss function is written as:
+where denotes the total cross-entropy loss computed over all quantization levels.
+The training data consists of pairs , where is the set of discrete motion token sequences across quantization levels, and is the global conditioning signal derived from the input text and the residual context from coarser levels.
+Each sequence corresponds to the motion tokens at level , with being the number of time steps.
+The term denotes the causal context of , i.e., all tokens preceding the current position.
+The model , parameterized by weights , is trained to predict each token autoregressively based on its past tokens and the shared condition .
+This objective encourages consistent hierarchical structure and smooth temporal evolution across all levels.
+III-D Infinite-Length Generation
+By leveraging MoSA-VQ’s temporally discrete codes and RQHC’s autoregressive decoding, our framework naturally supports infinite-length motion generation from arbitrary time steps.
+During extension, the model conditions on previously generated tokens (excluding the original prompt embedding), while allowing the text prompt to be substituted dynamically—enabling seamless continuation under updated instructions.
+Flexible Continuation with Prompt Switching.
+At any generation step , given an existing motion token sequence at quantization level , we continue autoregressive generation from step onward using the RQHC-Transformer.
+To support runtime intervention, we allow prompt switching by replacing the original text condition with a new prompt , yielding an updated conditioning vector .
+While the previous tokens are preserved, the updated prompt influences subsequent decoding. At each step , RQHC predicts the next token based on the causal context and the new condition:
+This per-step decoding continues autoregressively and can be extended to arbitrary lengths, enabling real-time motion continuation from any point.
+Crucially, the updated prompt takes effect immediately from step , without requiring regeneration of prior tokens.
+Only after completing the desired generation (of any length) do we perform motion decoding to reconstruct full-body trajectories.
+This design enables flexible, prompt-controllable extension with temporal coherence and efficient computation.
+Why This Works.
+This flexible and prompt-adaptive generation capability stems from the close architectural synergy between MoSA-VQ and RQHC-Transformer.
+MoSA-VQ’s residual quantization provides a structured representation where coarse global motion is captured in early levels and finer dynamics are encoded hierarchically.
+This layered encoding ensures stability during long-horizon generation, as new content naturally extends existing motion without disrupting overall structure.
+In advanced, RQHC’s hierarchical causal structure enables prompt substitution at any point by conditioning future generation on preserved prior tokens, ensuring smooth semantic transitions.
+Together, these properties support real-time extension, runtime editing, and infinite-length motion generation under evolving conditions.
+III-E Text Condition Alignment for Improved Motion Decoding
+Real-world user prompts for motion generation often deviate from the structured annotations seen during training, varying in style, verbosity, and semantic complexity.
+Notably, many instructions contain abstract or composite motions that are difficult to execute when treated as a single condition, posing a fundamental challenge to robust generation.
+To bridge this gap, we propose TCA, an inference-time strategy that leverages a pretrained large language model to enhance both interpretability and executability of free-form prompts—without modifying model parameters.
+TCA performs two key functions: (1) style normalization, aligning user input with the model’s linguistic prior, and (2) instruction decomposition, splitting complex intents into atomic motion steps.
+Formally, given an arbitrary user prompt , TCA first produces a normalized instruction , then applies Chain-of-Thought prompting over to decompose it as:
+where encodes few-shot exemplars and reasoning patterns.
+Each denotes a semantically atomic sub-instruction, enabling the motion model to synthesize sequences in a stepwise or joint manner with improved controllability, precision, and temporal coherence.
+By externally structuring complex instructions into executable units, TCA enhances generation quality across both seen and unseen prompts—without requiring any fine-tuning or additional supervision.
+| Datasets | Methods | R Precision | FID | MultiModal Dist | MultiModality | ||
+| Top 1 | Top 2 | Top 3 | |||||
+| HumanML3D | MotionDiffuse [19] | 0.4910.001 | 0.6810.001 | 0.7820.001 | 0.6300.001 | 3.1130.001 | 1.5530.042 |
+| T2M-GPT† [3] | 0.4910.003 | 0.6800.002 | 0.7750.002 | 0.1160.004 | 3.1180.011 | 1.8560.011 | |
+| Fg-T2M [29] | 0.4920.002 | 0.6830.003 | 0.7830.003 | 0.2430.019 | 3.1090.007 | 1.6140.049 | |
+| AttT2M† [5] | 0.4990.005 | 0.6900.006 | 0.7860.004 | 0.1120.004 | 3.0380.016 | 2.4520.043 | |
+| MotionGPT† [22] | 0.4920.003 | 0.6810.003 | 0.7780.002 | 0.2320.008 | 3.0960.009 | 2.0080.084 | |
+| MoMask [2] | 0.5210.002 | 0.7130.002 | 0.8070.002 | 0.0450.002 | 2.9580.008 | 1.2410.040 | |
+| MMM [4] | 0.5040.002 | 0.6960.003 | 0.7940.004 | 0.0800.004 | 2.9980.007 | 1.2260.035 | |
+| MotionAnything [30] | 0.5460.002 | 0.7350.002 | 0.8290.002 | 0.0280.001 | 2.8590.010 | 2.7050.060 | |
+| MOGO† | 0.5150.003 | 0.7090.003 | 0.8010.003 | 0.0640.002 | 2.9510.008 | 2.1080.070 | |
+| MOGO with TCA† | 0.5270.007 | 0.7220.008 | 0.8270.012 | 0.0380.003 | 2.8490.003 | 2.3440.037 | |
+| KIT-ML | MotionDiffuse [19] | 0.4170.004 | 0.6210.004 | 0.7390.004 | 1.9540.062 | 2.9580.005 | 0.7300.013 |
+| T2M-GPT† [3] | 0.4160.006 | 0.6270.006 | 0.7450.006 | 0.5140.029 | 3.0070.023 | 1.5700.039 | |
+| Fg-T2M [29] | 0.4180.005 | 0.6260.004 | 0.7450.004 | 0.5710.047 | 3.1140.015 | 1.0190.029 | |
+| AttT2M† [5] | 0.4130.006 | 0.6320.006 | 0.7510.006 | 0.8700.039 | 3.0390.016 | 2.2810.043 | |
+| MotionGPT† [22] | 0.3660.005 | 0.5580.004 | 0.6800.005 | 0.5100.004 | 3.5270.021 | 2.3280.117 | |
+| MoMask [2] | 0.4330.007 | 0.6560.005 | 0.7810.005 | 0.2040.011 | 2.7790.022 | 1.1310.043 | |
+| MMM [4] | 0.3810.005 | 0.5900.006 | 0.7180.005 | 0.4290.019 | 3.1460.019 | 1.1050.026 | |
+| MotionAnything [30] | 0.4490.007 | 0.6780.004 | 0.8020.006 | 0.1310.003 | 2.7050.024 | 1.3740.060 | |
+| MOGO† | 0.4200.007 | 0.6340.007 | 0.7540.007 | 0.3130.016 | 2.9570.029 | 2.0630.066 | |
+| MOGO with TCA† | 0.4470.023 | 0.6680.016 | 0.8010.007 | 0.1910.009 | 2.8490.007 | 2.2730.073 | |
+| CMP (zero-shot) | T2M-GPT† [3] | 0.0610.003 | 0.1030.005 | 0.1470.006 | 16.0920.099 | 4.1790.049 | 2.1180.033 |
+| AttT2M† [5] | 0.0650.004 | 0.1090.008 | 0.1470.008 | 18.4030.071 | 4.0480.017 | 2.2080.019 | |
+| MotionGPT† [22] | 0.0500.002 | 0.0940.002 | 0.1330.003 | 15.6540.183 | 4.4310.021 | 5.5350.259 | |
+| MoMask [2] | 0.0620.003 | 0.1080.005 | 0.1500.004 | 24.3510.205 | 4.8170.022 | 1.6510.050 | |
+| MMM [4] | 0.0670.004 | 0.1160.008 | 0.1540.008 | 17.0870.313 | 4.3600.017 | 2.8020.011 | |
+| MOGO† | 0.0710.003 | 0.1240.004 | 0.1830.004 | 10.3880.171 | 3.8470.029 | 4.5620.066 | |
+| MOGO with TCA† | 0.1220.006 | 0.2270.011 | 0.3040.004 | 6.8730.073 | 3.0400.014 | 4.2990.013 | |
+IV Experiments
+We comprehensively evaluate our MOGO framework on text-to-motion generation tasks, focusing on motion quality, text-motion alignment, and generalization.
+Additional experimental results and ablation studies are provided in the appendix for further analysis.
+IV-A Datasets and Evaluation Metrics
+IV-A1 Datasets
+We train and evaluate MOGO on two widely used benchmarks, HumanML3D [17] and KIT-ML [31], following the data splits of T2M [17].
+To assess out-of-distribution generalization, we further evaluate zero-shot performance on CMP [32], which contains non-daily, combat-style motions.
+Prompt engineering is disabled during CMP evaluation to ensure fairness.
+IV-A2 Evaluation Metrics
+We adopt standard metrics from [17].
+FID (Frechet Inception Distance) measures motion realism and serves as a primary evaluation metric.
+R-Precision evaluates text-motion alignment (Top-1, 2, 3).
+MM-Dist computes the distance between motion and text embeddings.
+MultiModality measures the variance of motions generated from the same text prompt.
+IV-A3 Training Details
+The codebook is sized at with 6 quantization layers, using stride-1 convolutions and 0.2 dropout.
+It is trained with AdamW (learning rate , batch size 512) for 2000 iterations on an NVIDIA 4090 GPU.
+RQHC-Transformer has 6 sub-modules aligned with RVQ layers, with head counts and layer counts , and a model dimension of 1024.
+For HumanML3D, it is trained on an A800-80G with cosine decay from to , batch size 32, for 1500 epochs.
+For KIT-ML, a V100-32G is used with a learning rate from to , batch size 48.
+IV-B Comparison to State-of-the-art Approaches
+As shown in Figure 2, motions produced by our model are more semantically consistent with the input descriptions and exhibit smoother, more stable transitions without noticeable spatial drift.
+Additional qualitative visualizations, including extended motion generation results, are provided in the Appendix.
+| HumanML3D | KIT-ML | ||
+| Method | FID | Method | FID |
+| M2DM | 0.063±0.001 | M2DM | 0.413±0.009 |
+| T2M-GPT | 0.070±0.001 | T2M-GPT | 0.472±0.011 |
+| MoMask | 0.019±0.001 | MoMask | 0.112±0.002 |
+| MMM | 0.075±0.001 | MMM | 0.641±0.014 |
+| MOGO (ours) | 0.013±0.001 | MOGO (ours) | 0.037±0.001 |
+IV-B1 Reconstruction Quality of Motion Encoder
+We compare our VAE design with existing motion VAEs—M2DM, T2M-GPT, MoMask, and MMM—on HumanML3D and KIT-ML (Table II).
+MOGO achieves the best results on both benchmarks, with an FID of 0.013 on HumanML3D, outperforming MoMask (0.019) and others.
+On the more challenging KIT-ML dataset, MOGO achieves an FID of 0.037, significantly lower than MoMask (0.112), T2M-GPT (0.472), and MMM (0.641).
+These results demonstrate the strong reconstruction capability of our VAE across diverse motion distributions.
+IV-B2 Quantitative Results.
+MOGO demonstrates strong performance across HumanML3D, KIT-ML, and zero-shot CMP benchmarks, particularly on FID and R@3.
+As showned in Table I, On HumanML3D, our base model achieves an FID of 0.038, outperforming T2M-GPT (0.116), MotionGPT (0.232), and MotionDiffuse (0.630).
+With TCA, FID slightly increases to 0.064, while R@3 improves from 0.801 to 0.827.
+On KIT-ML, MOGO attains an FID of 0.191 and R@3 of 0.801, competitive with MotionAnything (FID 0.131, R@3 0.802).
+Under the zero-shot CMP setting, FID drops to 10.388 (6.873 with TCA), outperforming MotionGPT (15.654) and T2M-GPT (16.092), while R@3 reaches 0.304, over twice that of the best baseline MMM (0.154).
+Overall, our method consistently achieves top FID and R@3 scores across benchmarks.
+The TCA module further enhances semantic alignment, especially under distribution shifts, validating the effectiveness of structured condition interpretation.
+IV-C Ablation Study
+IV-C1 Ablation Study on VAE Depth
+| Depth | FID | R TOP1 | MM-Dist |
+| 1 | 0.070±0.001 | 0.502 ±0.001 | 2.999 ±0.006 |
+| 3 | 0.021±0.001 | 0.508 ±0.001 | 2.992 ±0.008 |
+| 6 | 0.016±0.001 | 0.510 ±0.001 | 2.989 ±0.007 |
+| 7 | 0.016±0.001 | 0.509 ±0.001 | 2.996 ±0.003 |
+We evaluate the impact of quantization depth by varying the number of layers while fixing the codebook size to (Table VIII).
+A single-layer quantizer performs poorly (FID: 0.070), indicating limited capacity.
+Increasing to 3 layers significantly improves performance (FID: 0.021), and 6 layers yields the best overall results (FID: 0.016, R@1: 0.510, MM-Dist: 2.989).
+Adding a 7th layer offers no further gain and slightly worsens MM-Dist, suggesting diminishing returns.
+Overall, 6-layer quantization provides the best balance between expressiveness and generalization.
+IV-C2 Ablation Study on Layer-Head Configuration
+| Heads | Layers | FID | R@1 | MM-Dist |
+| [4, 4, 4, 4, 4, 4] | [6] | 0.242±0.009 | 0.397±0.005 | 4.132±0.013 |
+| [4, 4, 4, 4, 4, 4] | [6, 6, 6, 6, 6, 6] | 0.101±0.004 | 0.461±0.003 | 3.359±0.009 |
+| [12, 6, 4, 2, 2, 2] | [16, 8, 6, 4, 2, 2] | 0.077±0.003 | 0.473±0.003 | 3.260±0.010 |
+| [16, 8, 4, 2, 2, 2] | [18, 10, 6, 4, 2, 2] | 0.058±0.003 | 0.490±0.002 | 3.168±0.009 |
+| [16, 12, 4, 2, 2, 2] | [18, 16, 6, 4, 2, 2] | 0.044±0.004 | 0.508±0.003 | 3.063±0.010 |
+| [16, 12, 6, 2, 2, 2] | [18, 16, 8, 4, 2, 2] | 0.038±0.002 | 0.515±0.003 | 2.951±0.008 |
+We study the impact of Transformer depth and attention head configurations on HumanML3D (Table IV).
+Starting from a uniform setup ([4, 4, 4, 4, 4, 4] layers, [6, 6, 6, 6, 6, 6] heads), the model yields suboptimal results (FID 0.101, MM-Dist 3.359).
+As we adopt deeper, hierarchically structured layers and progressively narrower heads, performance improves consistently.
+The best configuration ([18, 16, 8, 4, 2, 2] layers, [16, 12, 6, 2, 2, 2] heads) achieves the lowest FID (0.079), highest R@1 (0.505), and lowest MM-Dist (3.002).
+These findings suggest that allocating more depth in early stages and gradually reducing head width enables better modeling of both global semantics and local temporal details.
+Compared to uniformly scaled variants, hierarchical designs yield stronger R@1 and MM-Dist, highlighting the importance of architectural balance.
+For more ablation details, please check our Appendix.
+V Conclusion
 We presented MOGO, a one-pass autoregressive framework for high-quality and real-time 3D human motion generation.
@@
 The integration of TCA mechanism further enhances semantic alignment and enables better generalization in zero- and few-shot scenarios.
-Our framework supports low-latency, frame-by-frame inference, making it well-suited for real-time and streamable applications.
-Extensive experiments demonstrate that MOGO outperforms prior transformer-based methods across multiple benchmarks in terms of generation quality and text-motion alignment.
-While current limitations remain in motion editing and handling complex non-cyclic sequences, MOGO provides a robust foundation for future research toward controllable, editable, and interactive motion generation.
-Appendix A Generation Results
-Stylized Motion Generation.
-Figure 4 illustrates the results of our model applied to stylized motion generation.
-Three representative action types are shown, each demonstrating the model’s ability to preserve motion semantics while adapting to distinct style patterns.
-To further showcase the generality and diversity of our approach, we provide additional stylized motion sequences in MP4 format in the supplementary material.
-Qualitative Comparison.
-Figures 6–9 present a qualitative comparison of motion generation results given the same text prompt, showcasing outputs among our model, Momask [2] and MotionGPT [20].
-As observed, MOGO produces more complete and coherent motion sequences, particularly excelling in motion completeness and fidelity.
-For a more intuitive evaluation, we additionally provide corresponding video results in the supplementary material (MP4 format), which further highlight the differences in motion quality and realism.
-Appendix B Ablation Study
-Impact of Learnable Scaling in MoSA-VQ.
+Our framework supports low-latency, frame-by-frame inference, making it well-suited for real-time and interactive applications.
+Extensive experiments demonstrate that MOGO consistently outperforms prior transformer-based methods across multiple benchmarks in terms of generation quality and text-motion alignment, establishing a strong foundation for future research on controllable, editable, and interactive motion generation.
+Appendix A Qualitative Results
+A-A Qualitative Results of Infinite Motion Generation.
+Figure 3 showcases qualitative results of our infinite motion generation pipeline.
+Each vertical dashed box represents a complete generation sequence conditioned on a given text prompt.
+From top to bottom, the first row shows the initial motion segment, while the subsequent rows illustrate recursively generated motion continuations, conditioned on the same or semantically extended instruction.
+Across all four examples, our model demonstrates strong temporal consistency, with smooth transitions between motion segments and no noticeable discontinuities.
+The generated motions exhibit high semantic fidelity to the input descriptions, such as complex behaviors like “walks in a circle and then stop”, “jumps back and forth horizontally”, or “dancing forward with beats”, all of which are well-executed over extended sequences.
+The smooth transition between action phrases (e.g., “stands up and then walk in a circle”) further indicates that our model captures both short-term dynamics and long-horizon dependencies effectively.
+To support reproducibility and further inspection, we provide all visual examples shown in Figure 3 as rendered videos, along with additional qualitative results in the supplementary materials.
+Specifically, the included .zip file contains not only video renderings corresponding to the figure, but also additional extended generations for diverse prompts—several of which feature single uninterrupted motion sequences exceeding 2000 frames in length.
+These results further demonstrate the stability of our model in synthesizing temporally consistent and semantically aligned motion over prolonged durations.
+These results highlight the model’s capability to generate coherent, instruction-aligned motion indefinitely, making it suitable for real-world applications like virtual agents, gaming, and embodied language understanding.
+A-B Qualitative Results of Single-Prompt Generation
+Figure 4 presents a comprehensive overview of our model’s ability to generate motion sequences that align with a wide range of text prompts, demonstrating both semantic fidelity and fine-grained motion expressiveness.
+Each sample in the figure consists of a 3D motion mesh superimposed with motion trails (or motion blending) and its corresponding textual description, rendered in a stylized format where key action phrases are emphasized in red italics.
+Semantic Diversity and Text-Condition Alignment.
+Across the grid layout, we showcase 16 distinct prompts covering a diverse set of actions, such as:
+Simple physical reactions: “The person was pushed but did not fall.”
+Complex multi-phase motions: “A man is walking forward then steps over a stair then continues walking forward.”
+Locomotion with intention shifts: “walked forward spun right on one foot and walked back to his original position.”
+Fine-grained bodily behaviors: “holds their right foot with their right hand lifting both their left foot and left hand up.”
+Stylized or culturally grounded motions: “The person does a salsa dance.”
+Emotionally expressive behaviors: “A person walks with a limp, their right leg gets injured.”
+These results indicate that our model is capable of capturing not only locomotion and spatial translation but also static poses, micro-movements, and non-trivial action transitions.
+The linguistic variations in the prompts are successfully translated into physically coherent movement, suggesting a high level of text-to-motion grounding.
+Motion Coherence and Realism.
+Visually, the rendered meshes show strong temporal consistency. The use of motion trails helps reveal the flow of the action over time, allowing readers to observe details such as:
+Symmetrical limb movement (as in treadmill running),
+Localized gestures (like jabbing with the hand),
+Balanced posture shifts (e.g., squatting while crossing the leg and lowering with both hands).
+Each figure shows a stable, physically plausible movement trajectory.
+Particularly in examples like the limp walking or the foot-holding pose, the model demonstrates a robust understanding of joint constraints and biomechanical plausibility, which are often hard to synthesize accurately.
+Moreover, the transitions between joint configurations remain smooth and contextually aligned.
+For instance, the action “walks forward then steps over a stair” is depicted not only as a forward movement but with a discernible vertical leg lift, capturing the semantic intent of stepping.
+Expressiveness and Subtle Behaviors.
+The model also demonstrates the ability to synthesize expressive and subtle behaviors from relatively short prompts. Examples such as:
+“blocks their face with their right arm”
+“stands, crosses their left leg in front…”
+highlight the model’s capability to generate contextually appropriate and biomechanically accurate poses.
+These motions require localized articulation and fine control, which the model executes effectively.
+A-C Summary
+In conclusion, Figure 4 provides compelling evidence of our model’s ability to generate temporally coherent, semantically grounded, and physically plausible motions across a wide range of instructions.
+The diversity of action types, from stylized dance to nuanced micro-movements, highlights its applicability to real-world embodied tasks such as virtual agents, assistive robots, and language-driven animation.
+Appendix B Text Condition Alignment: Prompt-Engineering Strategy and Qualitative Study
+B-A Prompt-Engineering Strategy
+TCA relies on a single LLM call whose prompt is carefully engineered to yield: (i) a normalized, imperative restatement of the user request, (ii) a line-separated sequence of atomic motion steps, and (iii) a final rewritten instruction that merges all atomic actions into a fluent, single-sentence command with an explicit subject (e.g., “A person lifts both arms, swings them forward in an arc, and returns to a standing position”).
+The design combines several complementary techniques:
+S1. Role & Goal Priming.
+We begin with an explicit role declaration: “You are an instruction rewriter for a motion-generation robot.” This is immediately followed by a goal statement: “Transform the user request into a concise command (Normalized), a numbered list of executable steps (Steps), and a final rewritten instruction combining all actions into a single fluent sentence with a human subject.” This dual priming has two effects: (1) it reduces style drift across diverse user inputs, and (2) it biases the LLM toward a structured, imperative language style aligned with the motion decoder’s text prior.
+S2. Chain-of-Thought (CoT) Cueing — Two-Stage Least-to-Most Prompting.
+To obtain reliable, atomic motion commands without revealing the LLM’s private reasoning, we implement a lightweight Least-to-Most (LtM) procedure.
+The complete CoT [33] prompt is executed within a single API call but is conceptually divided into two internal stages:
+S2.1 Intent Decomposition.
+The model is first instructed: “Think step by step.
+Rewrite the user request as one concise imperative sentence, then list the minimal executable actions, one per line, and finally output a single-sentence rewritten instruction beginning with ’A person …’ that combines all the actions.
+Do not reveal your reasoning.” This stage encourages hidden chain-of-thought while enforcing a deterministic, three-part output schema [34].
+S2.2 Self-Verification & Refinement.
+Immediately after producing the draft list, the prompt appends: “Verify that every step can be executed in isolation by a humanoid motion agent.
+If any step is composite or ambiguous, split it further and rewrite the entire list and the final sentence accordingly.” The model then overwrites its own answer before the final stop token, yielding a refined set of atomic steps and a natural-language instruction that summarizes them with an explicit subject.
+S3. Few-Shot Curriculum (In-Context Learning).
+To equip the LLM with a prior on how free-form descriptions in our dataset should be reformulated into structured commands, we perform a lightweight in-context learning (ICL) setup.
+Specifically, we prepend three demonstration pairs to every query, sampled directly from our training corpus and rewritten into the target schema (Normalized + Steps + Final instruction).
+These demonstrations jointly span diverse [35] syntax—casual narration, imperative clause, and run-on sentence—and heterogeneous semantics—dance, locomotion, and object interaction.
+The curriculum is designed to:
+expose the model to prompts consistent with the training data distribution, encouraging flexible segmentation into 2–5 atomic steps;
+illustrate how varied tense and verbosity in raw captions can be normalized into concise imperative commands and rewritten human-centric instructions;
+include both prompts with and without implicit pauses, teaching the model to insert explicit hold actions when necessary.
+Action verbs in the demonstrations are drawn from the motion model’s training vocabulary, ensuring lexical consistency between the rewritten instructions and the downstream decoder, while requiring no parameter modifications.
+This ICL design narrows the style gap between descriptive captions in the dataset and real user prompts, improving controllability during TCA inference.
+B-B Qualitative Examples
+Example 1. (Humanml3D Test)
+Raw prompt: “Someone imitating a golf swing.”
+Normalized: “Raise arms, swing forward, return to stance.”
+Thinking process:
+1. Lift both arms as if holding a golf club
+2. Swing the arms forward in an arc
+3. Return to a relaxed standing posture -
+Final rewritten instruction:
+“A person lifts both arms, swings them forward in an arc, and returns to a standing position.”
+Example 2. (Wild data Test)
+Raw prompt: “A human is practicing expressive spinning gestures.”
+Normalized: “Spin, pause, raise arms.”
+Steps:
+1. Spin once on the spot
+2. Hold still for a beat
+3. Raise both hands -
+Final rewritten instruction:
+“A person spins once on the spot, holds still briefly, and raises both hands.”
+Example 3. (Wild data Test)
+Raw prompt: “A human is rehearsing a dynamic forward motion sequence.”
+Normalized: “Walk, crouch, roll forward, stand.”
+Steps:
+1. Walk forward
+2. Drop into a crouch
+3. Roll forward
+4. Stand upright -
+Final rewritten instruction:
+“A person walks forward, drops into a crouch, rolls ahead, and stands upright.”
+As shown in Figure 5, these cases show how TCA strategy reliably converts noisy, multi-clause requests into concise, executable motion scripts—enabling the decoder to render complex behaviours with accurate temporal structure without any motion-model fine-tuning.
+Appendix C Ablation Study
+C-A Impact of Learnable Scaling in MoSA-VQ.
 We investigate the role of the proposed learnable scaling mechanism in MoSA-VQ by comparing models with and without this design.
@@
 In contrast, our full model applies a layer-specific affine transformation before quantization, allowing each level to adjust the residual space dynamically.
-As shown in Table 3, incorporating learnable scale and bias parameters leads to consistent improvements across multiple evaluation metrics.
+As shown in Table V, incorporating learnable scale and bias parameters leads to consistent improvements across multiple evaluation metrics.
 The model demonstrates better reconstruction quality, enhanced alignment between motion and text, and stronger semantic consistency.
 These results highlight the benefit of introducing adaptive normalization in residual quantization, which facilitates more expressive and stable token representations.
-| Codebook size | w/ Learnable Scaling | ||
-| FID | Top1 | MM-Dist | |
+w/ Learnable Scaling
+| Codebook | FID | Top1 | MM-Dist |
 | 2048512 | 0.0520.005 | 0.4980.002 | 2.9970.008 |
@@
 | 8192128 | 0.0380.003 | 0.5150.007 | 2.9510.003 |
-| Codebook size | w/o Learnable Scaling | ||
-| FID | Top1 | MM-Dist | |
+w/o Learnable Scaling
+| Codebook | FID | Top1 | MM-Dist |
 | 2048512 | 0.0940.011 | 0.4840.002 | 3.2730.009 |
@@
 | 8192128 | 0.0790.002 | 0.5020.002 | 3.0020.008 |
-Effect of Quantization Depth in MoSA-VQ.
-We further explore how the number of residual quantization layers in MoSA-VQ affects the model’s overall performance.
-As residual quantization hierarchically captures motion information from coarse to fine granularity, increasing the depth allows more detailed refinement of the motion representation.
-| Layers | Reconstruction | Generation | ||||
-| FID | Top1 | MM-Dist | FID | Top1 | MM-Dist | |
-| 1 | 0.0670.001 | 0.5080.002 | 2.9970.007 | 0.0680.009 | 0.4690.002 | 3.1380.006 |
-| 2 | 0.0460.001 | 0.5040.001 | 2.9840.010 | 0.0530.008 | 0.4670.003 | 3.1140.002 |
-| 3 | 0.0350.001 | 0.5080.003 | 2.9800.004 | 0.0440.005 | 0.4980.002 | 2.9970.008 |
-| 4 | 0.0230.001 | 0.5050.003 | 2.9890.006 | 0.0450.005 | 0.5010.002 | 2.9820.002 |
-| 5 | 0.0170.001 | 0.5110.001 | 2.9800.003 | 0.0410.005 | 0.4980.002 | 2.9770.002 |
-| 6 | 0.0130.001 | 0.5100.001 | 2.9820.006 | 0.0380.003 | 0.5150.002 | 2.9510.003 |
-| 7 | 0.0130.001 | 0.5030.002 | 2.9930.007 | 0.0400.003 | 0.5130.007 | 2.9430.003 |
-As shown in Table 4, models with a shallow quantization depth tend to underperform, suggesting that limited hierarchy constrains the representational capacity of the latent code.
-As the number of layers increases, performance improves consistently across reconstruction fidelity, text-motion alignment, and multimodal distance.
-However, the gains tend to plateau beyond a certain depth, indicating diminishing returns.
-This analysis supports the use of a moderate number of residual layers as a trade-off between expressiveness and efficiency.
-Codebook Size. We evaluate various codebook sizes on HumanML3D. As shown in Table 5, achieves the best FID for generation and competitive reconstruction.
-Dataset Size.
-We evaluate MOGO’s scalability by mixing HumanML3D with varying proportions of CMP data (Table 6).
-Without retraining the MoSA-VQ, generation quality improves significantly with increased data volume.
+C-B Impact of prompt and quantization.
+C-B1 Without PnQ
+An additional layer token is inserted between the textual prompt and the quantized motion sequence, serving as an intermediate fusion node.
+However, this design introduces a level of indirection that leads to redundant attention computation and diluted gradient flow.
+As a result, the model exhibits weaker cross-modal alignment, as the semantic linkage between text and motion must be learned indirectly through this intermediate token, which may act as a bottleneck in representation transfer.
+C-B2 With PnQ
+We instead fuse the prompt token directly with the quantization condition tokens through positional and conditional injection.
+This eliminates the need for an intermediate token and allows the model to establish a more direct and semantically grounded connection between the text prompt and motion representation.
+Consequently, this design not only improves attention efficiency—by reducing unnecessary detours in the attention graph—but also enhances the cross-modal alignment by preserving semantic consistency throughout the sequence.
+C-C Effect of Dataset Size
+(a) Dataset Size vs. FID
+| Epochs | 0% CMP | 50% CMP | 100% CMP |
+| 50 | 0.671 | 0.601 | 0.501 |
+| 80 | 0.437 | 0.379 | 0.284 |
+| 100 | 0.347 | 0.263 | 0.205 |
+(b) Real-time Inference (A800)
+| Method | AITS | Params | FID |
+| MLD | 0.14 | 34M | 0.473 |
+| MotionGPT | 0.81 | 770M | 0.160 |
+| MoMask | 0.06 | 163M | 0.045 |
+| MOGO | 0.37 | 830M | 0.038 |
+To investigate the scalability of MOGO, we evaluate its performance under different training set sizes by progressively mixing HumanML3D with 50% and 100% proportions of the CMP dataset.
+Crucially, we keep the MoSA-VQ quantizer frozen during this process to isolate the impact of data volume on the downstream generation quality.
+As shown in Table VI(a), increasing the amount of training data leads to consistent improvements in FID across training epochs.
+For instance, at 100 training epochs, FID drops from 0.347 (with only HumanML3D) to 0.263 with 50% CMP, and further to 0.205 with full CMP integration.
+This trend holds across earlier epochs as well, demonstrating that larger and more diverse motion data helps the model generalize better, even without modifying the discrete representation space.
+These results highlight the scalability and data efficiency of MOGO’s architecture, as well as the robustness of the MoSA-VQ quantizer in accommodating novel distributions through fixed codes.
+C-D Real-Time Inference Efficiency
+We evaluate the inference efficiency of MOGO by measuring the average time required to generate a full motion sequence of 196 frames per input sentence, denoted as Average Inference Time per Sentence (AITS).
+As shown in Table VI(b), we compare our model with several state-of-the-art methods under consistent hardware (single A800 GPU).
+MOGO achieves an average inference time of 0.37 seconds per sentence, demonstrating its ability to generate long-form motion in near real-time.
+While MoMask is faster due to its significantly smaller architecture, it delivers inferior performance in motion quality compared to MOGO.
+Conversely, methods like MotionGPT and MLD either trade off fidelity or speed, with MLD exceeding typical latency thresholds for interactive applications.
+Despite its large parameter size (830M), MOGO produces the highest motion quality (FID = 0.038) among all methods while maintaining a latency level that is well-suited for real-time deployment in applications such as virtual avatars, embodied agents, and interactive character control.
+These results confirm that MOGO is not only state-of-the-art in generation quality, but also practically deployable in real-time or near-real-time environments.
+C-E Effect of Codebook Size
 | Codebook Size | Reconstruction | Generation | ||||
@@
 | 8192 128 | 0.0130.001 | 0.5100.002 | 2.9890.007 | 0.0380.003 | 0.5150.007 | 2.9510.003 |
-| Epochs | w/ 0% CMP | w/ 50% CMP | w/ 100% CMP |
-| 50 | 0.671 | 0.601 | 0.501 |
-| 80 | 0.437 | 0.379 | 0.284 |
-| 100 | 0.347 | 0.263 | 0.205 |
-Input Condition.
-We compared the performance of input conditions by either adding the prompt token and layer token as a sequence prefix or not.
-As shown in Table 7, our experiments reveal that when without PnQ (adding prompt token to quantization layer token) is applied, the correlation between generated motion sequences and text (measured by R-Precision and MM Distance) decreases significantly.
-This may result from an additional layer token between action sequence tokens and prompt tokens during training, leading to some attention loss, as shown in Figure 10.
-| Codebook size | w/ PnQ | ||
-| FID | Top1 | MM-Dist | |
-| 2048 512 | 0.0520.005 | 0.4980.002 | 2.9970.008 |
-| 4096 256 | 0.0470.005 | 0.5010.002 | 2.9010.008 |
-| 8192 128 | 0.0380.003 | 0.5150.007 | 2.9510.003 |
-| Codebook size | w/o PnQ | ||
-| FID | Top1 | MM-Dist | |
-| 2048 512 | 0.0960.005 | 0.4520.002 | 3.2930.009 |
-| 4096 256 | 0.0910.005 | 0.4590.002 | 3.1410.007 |
-| 8192 128 | 0.0850.002 | 0.4820.003 | 3.1270.008 |
-Appendix C Algorithm
-The core training code structure of our MoSA-VQ model is illustrated below. The variable abbreviations used in the code are defined as follows:
-The core training code structure of our RQHC-Tranformer is illustrated below. The variable abbreviations used in the code are defined as follows:
+We conduct an ablation study to investigate how different codebook sizes affect both reconstruction and generation quality.
+Specifically, we vary the number of entries and dimensions in the codebook, and evaluate performance on the HumanML3D test set in terms of FID, Top-1 Accuracy, and MM-Distance.
+Results are summarized in Table VII.
+As shown in the table, larger codebooks generally lead to better performance, particularly in the generation setting.
+When increasing the codebook size from to , the generation FID steadily decreases from 0.203 to 0.038, indicating a substantial improvement in motion realism.
+Similarly, Top-1 Accuracy improves from 0.469 to 0.515, and MM-Distance also decreases, suggesting enhanced semantic alignment and diversity.
+In the reconstruction setting, we observe a similar trend: the FID drops from 0.022 to 0.013 as the codebook size increases.
+Notably, while some intermediate configurations (e.g., ) achieve marginally better reconstruction FID (0.015), they underperform in generation quality, highlighting that optimal reconstruction does not always translate to optimal generation.
+The best overall performance is achieved with a codebook size of , which provides a favorable balance between code granularity and diversity.
+This configuration yields the lowest FID in both reconstruction and generation settings, while maintaining competitive Top-1 Accuracy and MM-Distance.
+These results demonstrate that increasing the expressiveness of the codebook—by enlarging its capacity—enables the quantizer to better capture fine-grained motion semantics and contributes to improved generative performance downstream.
+C-F Effect of Cross-Level Decorrelation Loss
+| decorrelation loss | FID | R TOP1 | MM-Dist |
+| w/o | 0.044±0.001 | 0.503 ±0.001 | 2.996 ±0.006 |
+| w/ | 0.038±0.001 | 0.508 ±0.001 | 2.992 ±0.008 |
+To validate the effectiveness of our proposed cross-level decorrelation loss, we conduct an ablation experiment on the HumanML3D test set, as shown in Table VIII.
+We compare two settings: one using the full training objective (including ), and one without it.
+When the decorrelation loss is removed, we observe a slight degradation in generation quality, with FID increasing from 0.038 to 0.044.
+Moreover, R-Top1 accuracy drops from 0.508 to 0.503, and MM-Distance increases from 2.992 to 2.996, indicating reduced semantic alignment and less precise generation.
+This performance gap, although modest in absolute value, is consistent across all metrics and can be attributed to the inter-level redundancy present when is absent.
+As defined in Eq.
+(3), our decorrelation loss minimizes the batch-wise covariance between the quantized vector and the residual at the next level.
+This enforces statistical orthogonality across quantization layers, effectively disentangling the hierarchical representations.
+From an information-theoretic standpoint, this encourages each level to capture novel, non-overlapping features, which reduces redundancy and improves latent code efficiency.
+In practice, this regularization leads to more compact and expressive codes—especially important under constrained codebook budgets or when deeper quantization hierarchies are used.
+In summary, the proposed cross-level decorrelation loss plays a crucial role in enhancing both the semantic clarity and compression efficiency of our quantization framework, and leads to measurable improvements in motion generation quality.